40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 23 20:32:43 2015, Koji, Update, LSC, Frequency source fixed, IMC LO level adjusted 
    Reply  Mon Nov 23 21:48:49 2015, Koji, Update, LSC, Frequency source fixed, IMC LO level adjusted IMG_2136.JPGIMG_2129.JPGIMG_2133.JPG
       Reply  Thu Nov 26 19:39:27 2015, Koji, Update, LSC, Current state of the frequency source, and possible improvement RF_Frequency_Source.pdf
          Reply  Sun Nov 29 05:23:57 2015, rana, Update, LSC, Current state of the frequency source, and possible improvement 
             Reply  Sun Nov 29 12:32:26 2015, Koji, Update, LSC, Current state of the frequency source, and possible improvement 
Message ID: 11801     Entry time: Mon Nov 23 21:48:49 2015     In reply to: 11800     Reply to this: 11817
Author: Koji 
Type: Update 
Category: LSC 
Subject: Frequency source fixed, IMC LO level adjusted 

The trouble we had: the 29.5 MHz source had an output of 6 dBm instead of 13 dBm.

The cause of the issue: A short cable inside had its shield cut and had no connection of the return.


- The frequency source box was dismantled.
- The power supply voltages of +28 and +18 were provided from bench supplies.
- The 29.5 MHz output of 5~6 dBm was confirmed on the work bench.
- The 11 MHz OCXO out (unused) had an output of 13 dBm.

- Once the lid was opened, it was immediately found that the output cable for the 29.5 MHz source had a sharp cut of the shield (Attachment1).
- OK. This cable was replaced. The output of 13 dBm was recovered.

- But wait. Why is the decoupling capacitor on the 29.5 MHz OCXO bulging? The polarity of the electrolytic capacitor was wrong!
- OK. This capacitor was replaced. It was 100 uF 35 V but now it is 100 uF 50 V.

- I further found some cables which had flaky shields. Some of them were twisted. When the panel cable s connected, the feedthroughs were rotated. This twists internally connected cables. Solder balls were added to the connector to reinforce the cable end.

- When the box was dismantled, it was already noticed that some of the plastic screws to mount the internal copper heat sinks for ZHL-2's were broken.
They seemed to be degraded because of the silicone grease. I didn't try to replace all as it was expected to take too much time, so only the broken screws
were replaced with steel screws with shoulder washers
at the both side of the box.

- After confirming the circuit diagram, the box was returned to the rack. The 29.5 MHz output of 13 dBm there was confirmed.

Attachment 1: IMG_2136.JPG  542 kB  | Show | Hide all | Show all
Attachment 2: IMG_2129.JPG  944 kB  | Hide | Hide all | Show all
IMG_2129.JPG
Attachment 3: IMG_2133.JPG  5.984 MB  | Hide | Hide all | Show all
IMG_2133.JPG
ELOG V3.1.3-